# Are Large Language Models Memorizing Bug Benchmarks?

### v2

@@ II Methodology
We also collected a new dataset of 3214 GitHub repositories written in Java and Python. We targeted repositories from 2024 to reduce the likelihood that current state-of-the-art LLMs have seen them. To narrow our selection to likely high-quality, indicative repositories, we applied a minimum threshold of 100 stars, a proxy for community interest and engagement.
To ensure that we were mining novel datasets for LLMs, we applied the MinHash technique [16] with Locality Sensitive Hashing (LSH) [17] to filter repositories potentially overlapping with existing models’ training data. Because of code duplication, some files in new 2024 repositories might be identical to those in older repositories. As a result, we collected all repositories created between 2022 and the end of 2023 with greater than 100 stars to exclude the repositories from 2024 that contained data from these older, highly popular repositories. Finally, to manage computational resources, we randomly sampled files from the collected repositories ($250$ files per language) for our evaluation.
@@
We select a combination of models used for fault localization [25], program repair [26], and vulnerability detection [27]. Table II shows model information. [^1] We could not find detailed information about the training data for Mistral’s models, apart from anecdotal evidence on public forums. These models are from families of well-known code-related models, including codegen-multi, the LLaMa 3.1 family, the Gemma 2 family, StarCoder, and Mistral.
@@ III-B Negative Log Likelihood
Figure 3 shows NLL values for families of open-source models, allowing us to examine trends in familiarity across benchmarks. [^5] Trends are similar across models within a family; we provide full data at https://anonymous.4open.science/r/bugleakage-C786/. Note that Negative Log Likelihood (NLL) depends on tokenization and architecture. This means we can only directly compare NLL values within model families.
@@ III-D How Do Model Characteristics Influence Risk of Leakage?
Regression results reveal consistent trends across model families (e.g., LLaMa 3.1 or Gemma). For example, both LLaMa 3.1 8B and 70B were trained on the same data with a training budget of 15T tokens. However, the 70B model is approximately ten times larger than the 8B model, leading to an overall increase in 5-gram accuracy across all benchmarks (as shown in Figure 5) and decrease in NLL. Similarly, within the Gemma 2 family, the 2B model was trained on 3T tokens, while the 27B was trained on 12T. Here, we observe the same trend: average 5-gram accuracy increases in the 27B model and NLL decreases.
@@ V Limitations and Threats
Data Collection: Despite efforts to filter out older GitHub repositories when collecting new data, we anecdotally observed instances where files appeared to be adaptations of existing files (e.g., from 2018666https://bit.ly/3Od3nC6.) Our filtering process may not have perfectly excluded legacy code. We moreover cannot guarantee that the new repositories are identically distributed compared to those in the benchmarks. To mitigate this issue, given that the repositories in the benchmarks tend to be highly recognizable projects, we applied a >100-star filter to argue that the selected projects are of comparable quality.
@@ VII Conclusion
In this paper, we measure data leakage risks across multiple widely used bug benchmarks and state-of-the-art open-source models. Older benchmarks, especially Defects4J, exhibit higher memorization signals across all models, particularly in codegen-multi 6B. Newer datasets such as GitBug-Java, BugsInPy, BugsCpp show lower leakage risk, with similar NLL and 5-gram accuracy to new 2024 repositories. We also find that newer models with higher training budgets display significantly lower risks of data leakage. Our findings suggest that researchers should consider both the models they use in evaluation and pair older benchmarks, such as Defects4J, with more recent benchmarks, such as GitBug-Java, to improve model evaluation and address potential leakage issues appropriately.

### v3

@@ II Methodology
We collected a new dataset of 3,214 GitHub repositories written in Java (1,183) and Python (2,031), targeting repositories from 2024 to reduce the likelihood that state-of-the-art LLMs have seen them. To focus on high-quality repositories, we applied a 100-star minimum threshold as a proxy for community interest. To ensure novelty, we used MinHash [16] with Locality Sensitive Hashing (LSH) [17] to filter repositories overlapping with existing training data. Due to code duplication, some files in 2024 repositories might overlap with older repositories. Therefore, we included repositories from 2022–2023 (>100 stars) to exclude 2024 repositories containing duplicated data. Finally, we randomly sampled 250 files per language to manage computational resources for evaluation.
@@
We select a combination of models used for fault localization [25], program repair [26], and vulnerability detection [27]. Table II shows model information.These models are from families of well-known code-related models, including codegen-multi, LLaMa 3.1, Gemma 2, StarCoder, and Mistral.
@@ III-B Negative Log Likelihood
Figure 3 shows NLL values for families of open-source models, allowing us to examine trends in familiarity across benchmarks.Note that Negative Log Likelihood (NLL) depends on tokenization and architecture. This means we can only directly compare NLL values within model families.
@@ III-D How Do Model Characteristics Influence Risk of Leakage?
Regression results reveal consistent trends across model families. For example, both LLaMa 3.1 8B and 70B were trained on the same data with a training budget of 15T tokens. However, the 70B model is approximately ten times larger than the 8B model, leading to an overall increase in 5-gram accuracy across all benchmarks (as shown in Figure 5) and decrease in NLL. Similarly, within the Gemma 2 family, the 2B model was trained on 3T tokens, while the 27B was trained on 12T. Here, we observe the same trend: average 5-gram accuracy increases in the 27B model and NLL decreases.
@@ V Limitations and Threats
Data Collection: Despite efforts to filter out older GitHub repositories when collecting new data, we anecdotally observed instances where files appeared to be adaptations of existing files (e.g., from 2018) Our filtering process may not have perfectly excluded legacy code. We moreover cannot guarantee that the new repositories are identically distributed compared to those in the benchmarks. To mitigate this issue, given that the repositories in the benchmarks tend to be highly recognizable projects, we applied a >100-star filter to argue that the selected projects are of comparable quality.
@@ VII Conclusion
In this paper, we measure data leakage risks across multiple widely used bug benchmarks and state-of-the-art open-source models. Older benchmarks, especially Defects4J, exhibit higher memorization signals across all models, particularly in codegen-multi 6B. Newer datasets such as GitBug-Java, BugsInPy, BugsCpp show lower leakage risk, with similar NLL and 5-gram accuracy to new 2024 repositories. We also find that newer models with higher training budgets display significantly lower risks of data leakage. Our findings suggest that researchers should consider both the models they use in evaluation and pair older benchmarks, such as Defects4J, with more recent benchmarks, such as GitBug-Java, to improve model evaluation and address potential leakage issues
VIII Acknowledgments
The authors would like to thank Luke Dramko for providing code that we used for deduplicating 2024 repositories. This work was supported by Fundação para a Ciência e Tecnologia (FCT) through the CMU Portugal Dual PhD program: Daniel Ramos (SFRH/BD/150688/2020), Claudia Mamede (PRT/BD/155043/2022), Paulo Canelas (SFRH/BD/151469/2021), Catarina Gamboa (PRT/BD/154254/2021).